Finance question: Machine A has a value of 26 and will produce 30 in prosperity and can be abandoned for 19. Machine B has a value of 24 and will produce 30 in prosperity and 15 in a bad economy. Is Machine A worth 2 more? Please use a discount rate of 8%.
1
Expert's answer
2016-07-05T12:54:02-0400
If Machine A has a value of 26 and will produce 30 in prosperity and can be abandoned for 19, then with a discount rate of 8% the net present value of Machine A is:
NPV = -26 + 0.5*(30 + 19)/0.08 = $280.25.
If Machine B has a value of 24 and will produce 30 in prosperity and 15 in a bad economy, then with a discount rate of 8% the net present value of Machine B is:
NPV = -24 + 0.5*(30 + 15)/0.08 = $257.25. So, Machine A worth $23 more.